OpenStudy (anonymous):

find any values for which the rational expression is undefined 4x+1 / x^2+2x-3

OpenStudy (anonymous):

Same idea, x^2+2x-3=0 now factor it to find x

OpenStudy (anonymous):

3x^3=-3

OpenStudy (anonymous):

not sure where the power of 3 came from... x^2+2x-3 factors as (x-1)(x+3) so x=1 and x=-3

OpenStudy (anonymous):

thank you for the answer :)